OpenStudy (anonymous):

in base 10, two-digit prime n is 45 more than the number formed by the reversing the digits of n. find all possible values of n.

OpenStudy (anonymous):

Let \(n_1\) and \(n_2\) be the numbers in the 10s and 1s digit, respectively; then, \[n=10n_1+n_2\] The given info gives us the following equation: \[\begin{align*}10n_1+n_2&=45+10n_2+n_1\\ 10n_1+n_2&=10(n_2+4)+(n_1+5) \end{align*}\] Mathcing up the 10s and 1s digits of these numbers, you have \[\begin{cases}n_1=n_2+4\\n_2=n_1+5\end{cases}\] I'm afraid there aren't any solutions to this system, but it's possible my reasoning is flawed. I just thought I'd try my hand at it.

ganeshie8 (ganeshie8):

\[ \begin{cases}n_1=n_2+4\\n_2=n_1+5\end{cases} \] is valid only when \(n_1 \le 4\)

ganeshie8 (ganeshie8):

for \(n \gt 4\) we can use : \[ \begin{cases}n_1=n_2+5\\n_2=n_1-5\end{cases} \] \(\implies n_1 - n_2 = 5\) solving in single digit positive integers gives two solutions : \(61, 83\)